Andrea has an average of 87% on her first 5 tests. She has one test left to take. What is the lowest score Andrea can get and still have an average of at least 80% on her test?
A) 45 B) 73 C) 80 D) 43

Answer by checkley77(12569)

(87*5+x)/6>=80
(435+x)/6>=80
435+x>=6*80
435+x>=480
x>=480-435
x>=45 is the minimum test score to maintain an 80 average.
Proof:
(87*5+45)/6>=80
(435+45)/6>=80
480/6>=80
80>=80